OpenStudy (kirbykirby):

Does this sequence converge, if so, what's its limit? An = 1.01 A_n+1 = 1/2(An^2 + 1) for n>=1 Does this sequence converge, if so, what's its limit? An = 1.01 A_n+1 = 1/2(An^2 + 1) for n>=1 @Mathematics

OpenStudy (kirbykirby):

\[a _{n+1}=1/2((a _{n})^2+1)\]

OpenStudy (eyust707):

yes to 1

OpenStudy (eyust707):

well the way you have it written it would be yes to 0

OpenStudy (eyust707):

but im not really sure what you mean by an = 1.01

OpenStudy (kirbykirby):

Oops sorry A1 = 1.01

OpenStudy (kirbykirby):

how do we figure it out though

OpenStudy (eyust707):

what happens as n gets really really big?

OpenStudy (kirbykirby):

dont we need to prove tho that its bounded and increasing or decreasing tho?

OpenStudy (kirbykirby):

our prof said we need to do that for recursive sequences

OpenStudy (eyust707):

well its been a while since i did sequences but from what i remember you end up with a sequence of number like a1, a2, a3, a4, a5, ... an it converges if it exists and approaches some number

OpenStudy (eyust707):

in your case as n gets really big you get 1/infinity

OpenStudy (eyust707):

each number gets smaller and smaller until eventually it might as well be considered 0

OpenStudy (kirbykirby):

sorry its 1/2* (....) the An^2 part is in the numerator, only the "2" is in the denom

OpenStudy (eyust707):

in that case it would diverge off to infinity

OpenStudy (eyust707):

you can tell this by looking at what happens as you plug in n's

OpenStudy (eyust707):

the numbers get bigger and bigger

OpenStudy (kirbykirby):

but we have an example that's not completed, and it said that a similar sequence converged, and it was 1/2*(An+6) with A1=2

OpenStudy (kirbykirby):

even tho the first few terms were growing, it somehow converged

OpenStudy (eyust707):

hold on a sec i cant quite remember sequences

OpenStudy (kirbykirby):

ok

OpenStudy (eyust707):

lim as n approaches infinity in the sequence (1/2)(An+6) converges off to infinity

OpenStudy (eyust707):

i mean diverges off to infinity***

OpenStudy (kirbykirby):

I dunno I have that the limit is 6 in my book, but they dont really go into how they found it

OpenStudy (eyust707):

whats the question say exactly?

OpenStudy (kirbykirby):

t asks "Does the sequence converge? If so, find it's limit, Otherwise, say it diverges" and it shows the recursive sequence above

OpenStudy (eyust707):

i m 100 percent sure as you increase n in that situation the numbers get bigger than 6

OpenStudy (eyust707):

they must be asking for something different.. are you sure they dont give a bound on n?

OpenStudy (eyust707):

if n< or = 6 than it would converge to 6

OpenStudy (kirbykirby):

well they give n >= 1

OpenStudy (eyust707):

you or I must be missing something... not quite sure what. maybe one of the other guys in here could help ya i honestly dont know why that would = 6 and not infinity

OpenStudy (kirbykirby):

=\ ouille this is a toughy

OpenStudy (perl):

let me try

OpenStudy (perl):

the sequence converges, if lim An+1/An = 1 , i believe

OpenStudy (eyust707):

yes.. thats not what she gave tho

OpenStudy (eyust707):

she said (1/2)(An+6) ; n>1

OpenStudy (eyust707):

does the sequence converge, if so to what?

OpenStudy (perl):

if An is bounded and monotonic, you can use monotonic sequence theorem

OpenStudy (eyust707):

not sure what that is lol

OpenStudy (kirbykirby):

im not sure how to use it tho.. what do i show it's bounded above/below to?

OpenStudy (perl):

lets plug in some values to see where this is going

OpenStudy (perl):

heres another idea

OpenStudy (eyust707):

that thrm says that if An gets smaller and smaller than it has a limit

OpenStudy (perl):

solve the algebraic equation

OpenStudy (eyust707):

An is getting bigger

OpenStudy (kirbykirby):

when i ddo that i get like A1 = 1.01 A2 = 1.01005 A3 = 1.010100501 A4 = 1.010151511 A5 = 1.010203038

OpenStudy (perl):

A_n+1 = 1/2A_n^2 +1/2, now subtract both sides A_n+1

OpenStudy (eyust707):

ohh your looking at the top one

OpenStudy (perl):

0 = 1/2 A_n^2 - A_n+1 + 1/2 , multiply both sides by 2 now

OpenStudy (perl):

0 = A_n^2 - A_n+1 + 1

OpenStudy (perl):

since for large n , A_n ~ A_n+1 , we can substitute and make a quadratic

OpenStudy (perl):

err 0 = A_n^2 - 2A_n+1 + 1

OpenStudy (perl):

so first substitute A_n for A_n+1

OpenStudy (perl):

0 = A_n+1^2 - 2A_n+1 + 1 , 0 = ( A_n+1 -1) ( A_n+1 -1)

OpenStudy (perl):

A_n+1 = 1

OpenStudy (perl):

it is easier to see this if you plug in A1 = 2, then it goes to 1 quickly

OpenStudy (perl):

the limit is 1

OpenStudy (kirbykirby):

wow i never knew you could do that trick where A_n ~ A_n+1 o_O

OpenStudy (perl):

kirky, your values are wrong

OpenStudy (perl):

A1 = 1.01 A2 = 1.01005 A3 = 1.010100501 A4 = 1.010151511 A5 = 1.010203038 this is false

OpenStudy (perl):

A1 = 1.01 A2 = 1.005 A3 = 1.0025 A4 = 1.00125 A5= 1.000625

OpenStudy (perl):

wait i made a mistake

OpenStudy (perl):

wait, im wrong, this is divergent

OpenStudy (perl):

i made a mistake, this is only true if |A1| < 1

OpenStudy (kirbykirby):

oh

OpenStudy (kirbykirby):

is there a way to prove its divergent

OpenStudy (perl):

I think you can do a ratio test

OpenStudy (kirbykirby):

We didn't get to the point in class where were at ratio tests =\... And.. isnt that for series only anyways? o.O

OpenStudy (perl):

what i did only helps if we knew a priori that there was a limit

OpenStudy (kirbykirby):

hmm.. ok well I dunno I can't seem to figure this one out but thanks for ur help tho

OpenStudy (perl):

we can show that An+1 > An

OpenStudy (kirbykirby):

ya i tried that and it worked, but i don't know how to show it's bounded.. since I'm not sure which bound I'm supposed to look for?

OpenStudy (kirbykirby):

I don't know if it might be converging to a number or not o_o?

OpenStudy (perl):

i dont think its converging

OpenStudy (perl):

for instance try 1.5

OpenStudy (perl):

as A1, that blows up fast

OpenStudy (kirbykirby):

oh i see what u mean

OpenStudy (perl):

even 1.1 blows up pretty fast

OpenStudy (perl):

ok got it, it diverges

OpenStudy (perl):

I am absolutely sure

OpenStudy (perl):

because i sat there and hit that calculator key must be 100 times

OpenStudy (perl):

starting 1.01

OpenStudy (kirbykirby):

wow lol

OpenStudy (kirbykirby):

ok i'll trust this then... hehe thanks again for your help :)!!! really appreciated